OpenStudy (summersnow8):

A sealed tank at room temperature, 25 degrees Celsius, has 22.0 g of CO2. The gas has a pressure of 732 mm Hg. The tank is moved to a room kept at 12 degrees C and an additional 10.0 g of CO2 are added to the tank. What is the pressure (mm Hg) in the tank? The answer is supposedly 1.02 x 10^3 mm Hg

OpenStudy (summersnow8):

my work: T1 = 25 + 273 = 298 T2 = 12 + 273 = 285 P1 = 732 mm Hg P2 = ? n1 = mass / molar mass = 22.0 g / (12 + 16 *2) = 0.5 mol n2 = mass / molar mass = 10.0 g / (12 + 16 * 2) = 0.2272727 mol p2 = P1n2T2 / n1T1 p2 = 1540, but the answer is supposed to be 1020 mm Hg